OpenStudy (anonymous):

Write the equation of the line that passes through (–3, 4) with a slope of 2 in point-slope form.

OpenStudy (anonymous):

Standard point slope equation: y-y1=m(x-x1)

OpenStudy (anonymous):

plug in x1 & y1

OpenStudy (anonymous):

& plug in the slope which is '2'

OpenStudy (anonymous):

so you should have y-4=2(x--3)

OpenStudy (anonymous):

Thanks so much :)

OpenStudy (anonymous):

No problem :)